OpenStudy (anonymous):

2x - 6y = 5 x + y = 2 Solve the system of equations.

OpenStudy (anonymous):

multiply the 2nd eq by2 2x+2y=4 2x-6y=5 8y=-1 y= -1/8 find x

OpenStudy (anonymous):

If the second eqn is doubled and subtracted from first eqn, what u get?

OpenStudy (maheshmeghwal9):

If \[x+y=2\] then, \[y=-1/8\] therefore, by putting value of y in the eq. \[x+y=2\] u can also get the value of x:)